# A Computational Social Simulation of Ageing and Care Accessibility in Italian Inner Areas

### v2

@@ 4.3 System-level and micro-level effects
Figure 2: Comparison of ambulatory service locations in the two simulated configurations. Top: Scenario 1 (current distribution of ambulatory services) shows fragmented accessibility with Walkability Index values ranging from 12 to 47 (mean $\approx$ 27). Bottom: Scenario 2 (relocation to Villa Bernocchi) yields a more homogeneous spatial distribution and higher average walkability (mean $\approx$ 34), indicating improved accessibility for peripheral households while maintaining service coverage for central areas. Grey-scale points represent residential nodes weighted by walkability index (0-50). Source: OpenStreetMap [22], Simulation Data.
This divergence between macro-neutrality and micro-variation illustrates that spatial interventions reallocate relative advantage rather than uniformly improving welfare. Such heterogeneous effects, typical of complex adaptive systems, highlight the need for multi-scale evaluation when interpreting policy outcomes.
@@ 4.4 Sensitivity-driven insights
The sensitivity analysis clarifies mechanisms behind the scenario results. Spatial impedance parameters directly govern accessibility outcomes, while behavioural capacity parameters condition the translation of spatial gains into reduced burden. Consequently, spatial redesign alone may not yield measurable welfare improvements without parallel organisational or social-support measures (e.g., additional service days, mobility assistance, tele-care).
@@ 5 Conclusions and Policy Implications
Service relocation in Premeno produces aggregate neutrality with local redistribution: system-level indicators remain statistically unchanged, whereas micro-spatial metrics reveal differentiated gains and losses. Behavioural heterogeneity mediates these effects, indicating that infrastructure redesign must be complemented by organisational and social-support interventions to achieve equitable burden reduction. Beyond the case study, the framework is transferable to other European small-municipality contexts where ageing, morphology, and service distribution interact. Its modular architecture enables comparative analyses and scenario testing for evidence-based policy in health-access planning.
@@ Acknowledgments
Author contributions: R. Garrone: model design, implementation, experiments, analysis, writing. Please note that Professor Bandini provided supervision at an earlier stage of the research project but was not involved in the conception, writing, or approval of the present manuscript. Her name was initially included in error and has been removed accordingly.

### v3

@@ 4.3 System-level and micro-level effects
Figure 2: Comparison of ambulatory service locations in the two simulated configurations. Top: Scenario 1 (current distribution of ambulatory services) shows fragmented accessibility with Walkability Index values ranging from 12 to 47 (mean $\approx$ 27). Bottom: Scenario 2 (relocation to Villa Bernocchi) exhibits a more homogeneous spatial distribution and higher average walkability values (mean $\approx$ 34), relative to Scenario 1, while maintaining service coverage for central areas. Grey-scale points represent residential nodes weighted by Walkability Index (0–50). Values shown are node-level WKB scores, aggregated over runs as means. Source: OpenStreetMap [21], simulation data.
@@ 4.4 Sensitivity-driven insights
Beyond the specific case study, the results have implications for how ageing-care systems are conceptualised and analysed within simulation-based modelling and analysis. The model demonstrates how spatial morphology, behavioural constraints, and decentralised decision-making jointly shape emergent accessibility and burden patterns. The sensitivity analysis clarifies mechanisms behind the scenario results. Spatial impedance parameters directly govern accessibility outcomes, while behavioural capacity parameters condition the translation of spatial gains into reduced burden. Consequently, spatial redesign alone may not yield measurable welfare improvements without parallel organisational or social-support measures (e.g., additional service days, mobility assistance, tele-care). More generally, this study illustrates how agent-based simulation can be used to formalise structural trade-offs in ageing-care systems by linking micro-level behaviour, spatial structure, and emergent system-level outcomes, consistent with core principles of computational social science and complex adaptive systems theory [4, 14].
@@ 4.5 Model robustness and theoretical implications
4.6 Limitations and Future Work
Several limitations delimit the scope of the present work. Behavioural rules are stylised, institutional dynamics are exogenously specified, and health trajectories are simplified. The model does not capture endogenous policy adaptation, learning by institutions, or long-term demographic feedbacks. Accordingly, results should be interpreted as structural insights rather than empirical predictions.
Future research may extend the framework in three directions. First, incorporating adaptive policy agents would allow exploration of endogenous governance dynamics. Second, coupling the model with longitudinal empirical data could support calibration of behavioural rules and validation of emergent patterns. Third, comparative applications across multiple municipalities would enable systematic investigation of how morphology, demography, and service density jointly shape accessibility regimes.
@@ 5 Conclusions and Policy Implications
By making explicit the mechanisms through which spatial and behavioural factors interact, the study demonstrates how agent-based modelling can inform ageing and care policy by revealing not only whether interventions work on average, but for whom and where their consequences are most pronounced. Service relocation in Premeno produces aggregate neutrality with local redistribution: system-level indicators remain statistically unchanged, whereas micro-spatial metrics reveal differentiated gains and losses. Behavioural heterogeneity mediates these effects, indicating that infrastructure redesign must be complemented by organisational and social-support interventions to achieve equitable burden reduction. Beyond the case study, the framework is transferable to other European small-municipality contexts where ageing, morphology, and service distribution interact. Its modular architecture enables comparative analyses and scenario testing for evidence-based policy in health-access planning.